OpenStudy (anonymous):

Find the x- and y-intercepts of the graph of the equation? x^2+y^2=64 I changed it to slope-intercept which is y=x+64 right? 64 is the y-intercept right? what is the x intercept? and did I change the equation correctly?

OpenStudy (anonymous):

Ahhh, sorry, I am a bit confused so there are no x-intercepts? And what about y? :)

OpenStudy (unklerhaukus):

the two y intercepts are for when x=0 and the two x intercepts are when y=0

OpenStudy (anonymous):

So do I have to find a way to make x=0 and y=0? :)

OpenStudy (unklerhaukus):

when x =0 0^2+y^2 = 64 what can y be?

OpenStudy (anonymous):

Oh, 8? :)

OpenStudy (unklerhaukus):

yes 8 and ?

OpenStudy (anonymous):

-8? orrr? :)

OpenStudy (unklerhaukus):

good so the two y-intercepts are at (0,8), and (0,-8)

OpenStudy (unklerhaukus):

can you get the x-intercepts too?

OpenStudy (anonymous):

Okay, thank you very much! So, wouldn't x be the same? Or no? :)

OpenStudy (unklerhaukus):

the values will be the same, but the coordinate pairs will be the other way around

OpenStudy (anonymous):

Okay, I understand now! Thank you very much for your help! :)

OpenStudy (unklerhaukus):

no you have all four intercepts, can you predict the 2D shape that the relation represents ?

OpenStudy (anonymous):

hmmm........ is it a flat horizontal line?

OpenStudy (anonymous):

actually wait...

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

it's almost like a diamond??

OpenStudy (unklerhaukus):

it's a circle of radius 8! 

OpenStudy (anonymous):

Ohhhhhhhhhhhhhhhhhh, hahaha, I'm sorry, I'm bad at math. That makes sense, thank you for helping me so much! :)

OpenStudy (unklerhaukus):

you will remember the equation of a circle now (centred at the origin, with radius r) \[x^2 + y^2 = r^2\]

OpenStudy (anonymous):

Ohhh, I didn't even notice that, thank you, that will help me on my exam tomorrow :)